OpenStudy (anonymous):

Will Medal And Fan Which equations in point-slope form are equations of the line that passes through the point (4, 5) and (−3, −1)? Choose exactly two answers that are correct. A. y - 3 = 7/6 (x-1) B. y - 5 = 6/7 (x-4) C. y + 1 = 6/7 (x+3) D. y - 4 = 7/6 (x-5)

OpenStudy (anonymous):

@RainbowBubbles01

OpenStudy (anonymous):

@RainbowBubbles01

OpenStudy (anonymous):

lol ok lets see

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

https://www.desmos.com/calculator/vphpgabluz was having trouble with that

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

A. y - 3 = 7/6 (x-1) B. y - 5 = 6/7 (x-4) C. y + 1 = 6/7 (x+3) D. y - 4 = 7/6 (x-5)

OpenStudy (anonymous):

put those in each of those like y - 3 = 7/6 (x-1) put it into one of them then put onther one in them till one of them are on those fit

OpenStudy (anonymous):

lol wat

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

wait i think i know what ur saying

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

put the answer choices in the boxes on desmos? until it fits?

OpenStudy (anonymous):

yeah but put one in the first one the if it doesnt look like it is the answer click the X button next the the box and try the next one

OpenStudy (anonymous):

lol ok thx alot

OpenStudy (anonymous):

how old r u?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

ok so u can help me more often then

OpenStudy (anonymous):

13